OpenStudy (liziscuteaf):

13 - 2(1 - 15) ÷ 4 SIMPLIFY 4 ME PLZ

OpenStudy (3mar):

\[\huge 13-2(1-15)÷4\] \[\huge =13-2(-14)÷4\] \[\huge =13+28÷4\] \[\huge =13+(28÷4)=13+7=20\]